OpenStudy (anonymous):

elimination method x+8y=67 -x+3y=21

OpenStudy (anonymous):

x = 3 y = 8

OpenStudy (anonymous):

add both equation , it will eliminate x 11y=88 y=8 substitute y= 8 in any of the equation and find out the value of x -x+3y=21 -x=21-24 x=3